OpenStudy (anonymous):

y=e^(x^3-5x^2) , find y'

OpenStudy (alfie):

\[e^{x^3-5x^2}\] Chain rule: \[f'(x) = (3x^2-10x)e^{x^3-5x^2}\]

OpenStudy (alfie):

Steps are pretty simple. The derivative of e^blabla is e^blabla. The derivative of x^3 = 3x^2 and -5x^2 = -10x

OpenStudy (anonymous):

thanks!